OpenStudy (anonymous):

What is the slope of a line that passes through points (-2,1) and (4,4)?Choose from one of these answers:-1/2,1/3,1/2,2

OpenStudy (anonymous):

This should be in math.

OpenStudy (anonymous):

m= rise/run = y/x m = (1-4)/(-2-4) = (-3)/(-6) = 1/2 m = 1/2